State the slope and the y-intercept for the graph of y + 3x = -7
Answer by rfadrogane(214)

State the slope and the y-intercept for the graph of y + 3x = -7.
Sol'n:
from the slope-intercept form:
y = mx + b, where m - slope, b - y-intercept
so, y + 3x = -7
y = -3x - 7
and finally, m = -3 & b = -7 --- answer